# Coherence and Transients in Nonlocally Coupled Dissipative Kicked Rotors

### v1

@@ Appendix A Coherent regions for coupled Chebyshev maps
The $n$th-order ($n\geq 2$) Chebyshev maps [31] are defined via $T_{n}(x)=\cos(n\arccos(x))$, $x\in[-1,1]$, which can be written as polynomials, for example, $T_{2}(x)=2x^{2}-1$, $T_{3}(x)=4x^{3}-3x$, $T_{4}(x)=8x^{4}-8x^{2}+1$, and so on. The $2$nd-order Chebyshev map is equivalent to the logistic map, except that it is upside-down and defined on the interval $[-1,1]$. Therefore, the dynamics of the coupled $T_{2}$ system is expected to resemble that of the coupled chaotic logistic map [21], as confirmed in the first row of Fig.7. For higher orders $n$, the phase diagrams show similar patterns but the sizes of the tongue regions vary; see the last two rows in Fig.7 for $n=3,4$. In each left panel, the yellow region indicates a fully synchronized state, where the temporal dynamics follows the single Chebyshev map, which is chaotic.
@@
Figure 7: Coherent regions for the system of nonlocally coupled Chebyshev maps $T_{n}$: $T_{2}$ (row 1), $T_{3}$ (row 2) and $T_{4}$ (row 3). The coupling scheme is the same as in Eq.(1). The left column shows wavenumbers $l=1,2,3$ with temporal period $\tau\leq 8$ in green, brown and red, respectively, and the fully synchronized region in yellow; the right column shows temporal periods $\tau=2,4,8$ in magenta, light blue and deep blue, respectively. All other patterns are colored in gray. Numerical settings: system size $N=300$, $P_{c}\times K\in[1,149]\times[0.01,1]$ of resolution $149\times 100$.
@@
Appendix B Coherent regions for coupled kicked rotors of larger system sizes
@@
Figure 8: Regions of typical wave numbers $l$ and temporal periods $\tau$ in the parameter $(P_{c},K)$-plane for the system sizes $N=300$ (row 1) and $1000$ (row 2). $P_{c}\in\{1,...,\frac{N}{2}-1\}$ and $K\in[0.01,5]$ for $100$ equidistant values. Each pixel color is decided by the mode of $10$ trajectories. Other parameter values are as in Fig.3.

### v2

@@ Appendix C Coherent regions for coupled Chebyshev maps
The $n$th-order ($n\geq 2$) Chebyshev maps [36] are defined via $T_{n}(x)=\cos(n\arccos(x))$, $x\in[-1,1]$, which can be written as polynomials, for example, $T_{2}(x)=2x^{2}-1$, $T_{3}(x)=4x^{3}-3x$, $T_{4}(x)=8x^{4}-8x^{2}+1$, and so on. The 22nd-order Chebyshev map is equivalent to the logistic map, except that it is upside-down and defined on the interval [1,1][-1,1]. Therefore, the dynamics of the coupled $T_{2}$ system is expected to resemble that of the coupled chaotic logistic map [25], as confirmed in the first row of Fig.A6. For higher orders $n$, the phase diagrams show similar patterns but the sizes of the tongue regions vary; see the last two rows in Fig.A6 for $n=3,4$. In each left panel, the yellow region indicates a fully synchronized state, where the temporal dynamics follows the single Chebyshev map, which is chaotic.
@@
Figure A6: Coherent regions for the system of nonlocally coupled Chebyshev maps $T_{n}$: (a)-(b) $n=2$, (c)-(d) $n=3$ and (e)-(f) $n=4$. The coupling scheme is the same as in Eq.(1). The left column shows wavenumbers $l=1,2,3$ with temporal period $\tau\leq 8$ in green, brown and red, respectively, and the fully synchronized region in yellow; the right column shows temporal periods $\tau=2,4,8$ in magenta, light blue and deep blue, respectively. All other patterns are colored in gray. Numerical settings: system size $N=300$, $P_{c}\times K\in[1,149]\times[0.01,1]$ of resolution $149\times 100$.
@@
Appendix D Coherent regions for coupled kicked rotors of larger system sizes
@@
Figure A7: Regions of typical wave numbers $l$ and temporal periods $\tau$ in the parameter $(P_{c},K)$-plane for the system size (a)-(b) $N=300$ and (c)-(d) $1000$. $P_{c}\in\{1,...,\frac{N}{2}-1\}$ and $K\in[0.01,5]$ for $100$ equidistant values. Each pixel color is decided by the mode over $10$ trajectories. Other parameter values are as in Fig.3.